OpenStudy (anonymous):

Consider both Earth and Mars to be spheres. The surface area of mars is about 28% that of earth. What is the ratio of their radii?

OpenStudy (anonymous):

surface area of a sphere = 4 pi r^2 if radis earth = re and of mars is rm then 4 pi rm^2 28 -------- = --- 4 pi re^2 100 can you continue from here?

OpenStudy (anonymous):

from the above rm^2 / re^2 = 28 / 100 so rm / re = sqrt28 / 10

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, you need any more help?

OpenStudy (anonymous):

how did you do your simplifying? My book got 1.89

OpenStudy (anonymous):

hmmm sqrt28 = 5. 29 so ratio is 0.53 to 2 d p 0h - ratio of mars to earth = 100 / 0.53 = 1 : 1.89

OpenStudy (anonymous):

or put another way the radius of the earth = 1.89 times the radius of mars

OpenStudy (anonymous):

i worked out radius of mars as fraction of the earth's hope thats clear